# The separating Noether number of small groups

### v1

@@ Abstract.
Partially supported by the Hungarian National Research, Development and Innovation Office, NKFIH K 138828.
@@ 1. Introduction
It is well known that for $v,w\in V$ with $G\cdot v\neq G\cdot w$, there exists an invariant $f\in K[V]^{G}$ with $f(v)\neq f(w)$ (see for example [16, Theorem 3.12.1], [33, Theorem 16] or [23, Lemma 3.1]). Therefore the general notion of a ”separating set of invariants” (see [16, Definition 2.4.1]) reduces to the following in the case of finite groups:
@@
The result mentioned before Definition 1.1 implies that every generating set of the algebra $K[V]^{G}$ is a separating set. In the structural study of algebras of polynomial invariants of finite groups the focus on separating sets (instead of the more special generating sets) turned out to be useful in finding strengthened versions of known theorems in the invariant theory of finite groups, see for example [24], [25]. Considering explicit computations of invariants, the class of modular representations for which generators of the corresponding algebra of invariants is known is very limited. On the other hand, in several such cases explicit finite separating sets are known, see for example [48], [37]. In the non-modular case, an explicit separating set of multisymmetric polynomials is given in [39]. The systematic study of separating invariants over finite fields began recently in [34], [38].
@@ 1.1. Degree bounds
Following [36, Definition 2] (see also [33]) we set
@@
A major difference between the Noether number and its separating version is that if the characteristic of the base field $K$ divides the order of $G$, then $\beta^{K}(G)=\infty$ by [42], whereas $\beta_{\mathrm{sep}}^{K}(G)\leq|G|$ also in the modular case (see for example the proof of [16, Theorem 3.12.1], [33, Theorem 16] or [23, Lemma 3.1]) However, in the present paper our focus is on the non-modular setup, when the characteristic of the base field does not divide the order of $G$. In this context an interesting and generally open question is the following:
@@
The exact value of $\beta_{\mathrm{sep}}^{K}(G)$ for certain finite abelian groups $G$ and an algebraically closed base field of non-modular characteristic was determined recently in [44], [45]. As far as we know, the only non-abelian finite groups whose separating Noether number in the non-modular case is determined in the literature are the non-abelian groups of order $3p$ for some odd prime $p$ (see [9]). For some results on the modular case see [36, Theorem C], [26].
@@ 2.1. Failure of strict monotonicity
$\beta^{K}(G)\geq\beta^{K}(G/N)\text{ and }\beta_{\mathrm{sep}}^{K}(G)\geq\beta_{\mathrm{sep}}^{K}(G/N).$
@@
(see [36, Theorem B]). The inequality (2) was sharpened in [13, Theorem 1.2], where the following strict monotonicity of the Noether number was proved:
@@ 3.1. Dependence on the base field of $\beta_{\mathrm{sep}}^{K}(G)$
(i) First we mention an example in the modular case. Consider the $\mathrm{S}_{n}$-module $\mathbb{F}_{q}^{n}$ endowed with the natural permutation representation, where $q$ is a prime power and $\mathbb{F}_{q}$ is the finite field with $q$ elements. Separating $\mathrm{S}_{n}$-invariants on $\mathbb{F}_{q}^{n}$ were studied (much before the notion of ’separating invariants’ was formally introduced) in [27], [1], and recently in [34], [20]. In particular, we have $\beta_{\mathrm{sep}}(\mathrm{S}_{3},\mathbb{F}_{2}^{3})=2$ by [34, Lemma 4.3] and $\beta_{\mathrm{sep}}(\mathrm{S}_{3},\mathbb{F}_{4}^{3})=3$ by [20, Corollary 4.9].
(ii) An example where the inequality of Lemma 3.1 (ii) is strict and $\mathrm{char}(K)$ does not divide $|G|$ will be provided in the present paper in Proposition 5.6. In the subsequent paper [21] we shall give an example where the inequality of Lemma 3.1 (iii) is strict and $\mathrm{char}(K)$ does not divide $|G|$.
@@
In contrast with Remark 3.2, it is shown in [35, Corollary 4.2] that $\beta^{K}(G)$ depends only on $\mathrm{char}(K)$, so
@@
It is interesting to compare this with (5), where the inequality for the corresponding separating Noether numbers is proved in the reverse direction. Of course, it maybe that both (5) and (6) always hold with equality. At least no example is known for $p$ and $G$ (where $p\nmid|G|$) for which any of (5) or (6) is a strict inequality. In fact (6) holds with equality for all $p\nmid|G|$ when $G$ is abelian, or when $G$ is one of the groups for which the exact value of their Noether number is given in [15]. For sake of completeness of the picture, we repeat that $\beta^{\mathbb{F}_{p}}(G)=\infty$ if $p$ divides $|G|$ by [42], whereas $\beta_{\mathrm{sep}}^{K}(G)\leq|G|$ for any $K$ (including the case when $\mathrm{char}(K)$ divides $|G|$).
@@ 3.3. Direct sum decompositions and multigradings
View the dual space $V_{i}^{*}$ as the subspace of $V^{*}$ consisting of the linear forms vanishing on $V_{j}$ for $j\neq i$. Then $V^{*}=\bigoplus_{i=1}^{k}V_{i}^{*}$. Choose a dual basis $x_{i}^{(j)}$ ($i=1,\dots,\dim(V_{j})$) in $V_{i}^{*}$, then our set of variables in $K[V]$ is $\{x_{i}^{(j)}\mid\ j=1,\dots,k;\ i=1,\dots,\dim(V_{j})\}$. Associated to the decomposition (7) there is an $\mathbb{N}_{0}^{k}$-grading on $K[V]$: the multihomogeneous component of multidegree $\alpha=(\alpha_{1},\dots,\alpha_{k})$ in $K[V]$ is spanned by the monomials having degree $\alpha_{j}$ in the variables belonging to $V_{j}^{*}$ for each $j=1,\dots,k$. Clearly, the $G$-action on $K[V]$ preserves this multigrading, and thus $K[V]^{G}$ is spanned by multihomogeneous elements. In the sequel whenever we mention a ‘multigrading’ not explicitly defined, this should refer to the multigrading that comes from a direct sum decomposition as in (7), which should be clear from the context. A typical way by which we shall exploit the multigrading is captured by the following obvious statement:
@@ 3.7. Relative invariants and the Hilbert ideal
For example, if $V$ is as in (10), then the variable $t_{\chi}$ is a relative invariant of weight $\chi$. Moreover, for a sequence $\chi^{(1)},\dots,\chi^{(k)}$ of elements of $\widehat{G}$, $t_{\chi^{(1)}}\cdots t_{\chi^{(k)}}$ is an indecomposable $G$-invariant if and only if $\chi^{(1)}\cdots\chi^{(k)}=1\in\widehat{G}$ and there is no $\ell<k$, $1\leq i_{1}<\cdots<i_{\ell}\leq k$ with $\chi^{(i_{1})}\cdots\chi^{(i_{\ell})}=1\in\widehat{G}$ (we say in this case that $\chi^{(1)},\dots,\chi^{(k)}$ is an irreducible product-one sequence over $\widehat{G}$, and we call $k$ its length). We say that $\chi^{(1)},\dots,\chi^{(k)}$ is a product-one free sequence over $\widehat{G}$ if there is no $1\leq i_{1}<\dots<i_{\ell}\leq k$ with $\chi^{(i_{1})}\cdots\chi^{(i_{\ell})}=1\in\widehat{G}$. The Davenport constant $\mathsf{D}(\widehat{G})$ of the finite abelian group $G$ is defined as the maximal length of an irreducible product-one sequence over $G$. Clearly, the maximal possible length of a product-one free sequence is $\mathsf{D}(\widehat{G})-1$.
@@ 4. Groups with a cyclic subgroup of index two
Note that our representation is the direct sum of a $2$-dimensional and a $1$-dimensional representation. So by Lemma 3.5 it is sufficient to show that $v$ and $v^{\prime}$ can not be separated by a multihomogeneous invariant (cf. Section 3.3) of degree at most $m$. Suppose for contradiction that $f(v)\neq f(v^{\prime})$, where $f$ is a multihomogeneous invariant of degree at most $m$. So $f=t^{d}f_{1}(x_{1},x_{2})$. Since the projections of these two vectors in the $x_{1},x_{2}$ coordinate plane are the same, no invariant depending only on $x_{1}$ and $x_{2}$ can separate them. Thus we have $d>0$. The invariants depending only on $t$ are the polynomials in $t^{2}$, which takes the same value (namely $1$) on these vectors, hence $\deg(f_{1})>0$. Since $t$ is $A$-invariant, $f_{1}(x_{1},x_{2})\in K[x_{1},x_{2}]^{A}$, so $f_{1}$ is a $K$-linear combination of $A$-invariant monomials in $x_{1}$ and $x_{2}$. The monomials having positive degree in $x_{2}$ vanish on both $v$ and $v^{\prime}$. It follows that $f_{1}$ must have an $A$-invariant monomial depending only on $x_{1}$. The only such monomials are the powers of $x_{1}^{m}$. Thus $\deg(f)\geq m+d>m$, a contradiction.
@@ 4.1. The groups $\mathrm{D}_{2n}\times\mathrm{C}_{2}$
implying in turn that $\beta_{\mathrm{sep}}^{K}(\mathrm{D}_{2n}\times\mathrm{C}_{2})\leq n+2$. The reverse inequality is given in Proposition 4.3. $\square$
@@ 5.2. The binary tetrahedral group
In this section $G:=\widetilde{\mathrm{A}}_{4}$ is the binary tetrahedral group: the special orthogonal group $\mathrm{SO}(3,\mathbb{R})$ has a subgroup isomorphic to $\mathrm{A}_{4}$ (the group of rotations mapping a regular tetrahedron into itself), and $G$ is isomorphic to the preimage of $\mathrm{A}_{4}$ under the classical two-fold covering $\mathrm{SU}(2,\mathbb{C})\to\mathrm{SO}(3,\mathbb{R})$. So $G$ can be embedded as a subgroup into the special unitary group $\mathrm{SU}(2,\mathbb{C})$, acting via its defining representation on $V:=\mathbb{C}^{2}$. Thus $V$ is a $G$-module, and it is well known that $\mathbb{C}[V]^{G}$ is generated by three homogeneous invariants $f_{6}$, $f_{8}$, $f_{12}$, having degree $6$, $8$, and $12$ (see [31, Lemma 4.1] or [41, Section 0.13]). Since $G$ (as a subgroup of $\mathrm{GL}(V)$) is not generated by pseudo-reflections, $\mathbb{C}[V]^{G}$ does not contain an algebraically independent separating set by [24], so the invariants $f_{6}$ and $f_{8}$ can not separate all the $G$-orbits in $V$. This implies that $\beta_{\mathrm{sep}}^{\mathbb{C}}(G,V)\geq 12$. A second proof valid for more general base fields is given below.
@@
having degree $4,6,12,12$. In fact $h_{12}^{(1)}$ can be omitted, because we have $h_{12}^{(1)}=-4h_{6}^{2}$. Set $v:=[1,\xi^{2}]^{T}$ and $v^{\prime}:=\xi v$. Then $(x_{1}^{2}+x_{2}^{2})(v)=0=(x_{1}^{2}+x_{2}^{2})(v^{\prime})$, hence $f_{2}(v)=0=f_{2}(v^{\prime})$. It follows that $h_{6}(v)=0=h_{6}(v^{\prime})$. Moreover, $h_{8}(v)=16(\xi^{4})^{2}=16=16(\xi^{8})^{2}=h_{8}(v^{\prime})$. Thus no invariants of degree less than $12$ separate $v$ and $v^{\prime}$. However, they have different $G$-orbit, since $h_{12}^{(2)}(v)=4^{3}(\xi^{4})^{3}=-64$, whereas $h_{12}^{(2)}(v^{\prime})=4^{3}(\xi^{8})^{3}=64$. Thus we proved the inequality $\beta_{\mathrm{sep}}(G,V)\geq 12$, implying in turn $\beta_{\mathrm{sep}}^{K}(G)\geq 12$. On the other hand, by [11, Corollary 3.6] we know that $\beta^{K}(G)=12$, implying the reverse inequality $\beta_{\mathrm{sep}}^{K}(G)\leq 12$. ∎
@@ 5.6. The non-abelian group $(\mathrm{C}_{2}\times\mathrm{C}_{2})\rtimes\mathrm{C}_{4}$
has $c^{2}$ in its kernel and gives the defining representation of $\mathrm{D}_{8}$ as the group of symmetries of a square in the euclidean plane. The element $ab$ also belongs to the center of $G$, and the factor group $G/\langle ab\rangle\cong\mathrm{C}_{4}\times\mathrm{C}_{2}$, generated by the coset of $c$ and the coset of $a$. Thus $G$ has the $1$-dimensional representations

### v2

@@ 1. Introduction
It is well known that for $v,w\in V$ with $G\cdot v\neq G\cdot w$, there exists an invariant $f\in K[V]^{G}$ with $f(v)\neq f(w)$ (see for example [16, Theorem 3.12.1], [34, Theorem 16] or [23, Lemma 3.1]). Therefore the general notion of a ”separating set of invariants” (see [16, Definition 2.4.1]) reduces to the following in the case of finite groups:
@@
The result mentioned before Definition 1.1 implies that every generating set of the algebra $K[V]^{G}$ is a separating set. In the structural study of algebras of polynomial invariants of finite groups the focus on separating sets (instead of the more special generating sets) turned out to be useful in finding strengthened versions of known theorems in the invariant theory of finite groups, see for example [24], [25]. Considering explicit computations of invariants, the class of modular representations for which generators of the corresponding algebra of invariants is known is very limited. On the other hand, in several such cases explicit finite separating sets are known, see for example [49], [38]. The minimal possible size of a separating set is investigated in [26]. In the non-modular case, an explicit separating set of multisymmetric polynomials is given in [40]. The systematic study of separating invariants over finite fields began recently in [35], [39].
@@ 1.1. Degree bounds
Following [37, Definition 2] (see also [34]) we set
@@
A major difference between the Noether number and its separating version is that if the characteristic of the base field $K$ divides the order of $G$, then $\beta^{K}(G)=\infty$ by [43], whereas $\beta_{\mathrm{sep}}^{K}(G)\leq|G|$ also in the modular case (see for example the proof of [16, Theorem 3.12.1], [34, Theorem 16] or [23, Lemma 3.1]) However, in the present paper our focus is on the non-modular setup, when the characteristic of the base field does not divide the order of $G$. In this context an interesting and generally open question is the following:
@@
The exact value of $\beta_{\mathrm{sep}}^{K}(G)$ for certain finite abelian groups $G$ and an algebraically closed base field of non-modular characteristic was determined recently in [45], [46]. As far as we know, the only non-abelian finite groups whose separating Noether number in the non-modular case is determined in the literature are the non-abelian groups of order $3p$ for some prime $p$ congruent to $1$ modulo $3$ (see [9]). For some results on the modular case see [37, Theorem C], [27].
@@ 2.1. Failure of strict monotonicity
(2) $\beta^{K}(G)\geq\beta^{K}(G/N)\text{ and }\beta_{\mathrm{sep}}^{K}(G)\geq\beta_{\mathrm{sep}}^{K}(G/N).$
@@
(see [37, Theorem B]). The inequality (3) was sharpened in [13, Theorem 1.2], where the following strict monotonicity of the Noether number was proved:
@@ 3.1. Dependence on the base field of $\beta_{\mathrm{sep}}^{K}(G)$
(i) First we mention an example in the modular case. Consider the $\mathrm{S}_{n}$-module $\mathbb{F}_{q}^{n}$ endowed with the natural permutation representation, where $q$ is a prime power and $\mathbb{F}_{q}$ is the finite field with $q$ elements. Separating $\mathrm{S}_{n}$-invariants on $\mathbb{F}_{q}^{n}$ were studied (much before the notion of ’separating invariants’ was formally introduced) in [28], [1], and recently in [35], [20]. In particular, we have $\beta_{\mathrm{sep}}(\mathrm{S}_{3},\mathbb{F}_{2}^{3})=2$ by [35, Lemma 4.3] and $\beta_{\mathrm{sep}}(\mathrm{S}_{3},\mathbb{F}_{4}^{3})=3$ by [20, Corollary 4.9].
(ii) An example where the inequality of Lemma 3.1 (ii) is strict and $\mathrm{char}(K)$ does not divide $|G|$ will be provided in the present paper in Proposition 5.6. In the subsequent paper [21] we shall give examples where the inequality of Lemma 3.1 (iii) is strict and $\mathrm{char}(K)$ does not divide $|G|$.
@@
In contrast with Remark 3.2, it is shown in [36, Corollary 4.2] that $\beta^{K}(G)$ depends only on $\mathrm{char}(K)$, so
@@
It is interesting to compare this with (6), where the inequality for the corresponding separating Noether numbers is proved in the reverse direction. We mention that no example is known for $p$ and $G$ (where $p\nmid|G|$) for which (7) is a strict inequality. In fact (7) holds with equality for all $p\nmid|G|$ when $G$ is abelian, or when $G$ is one of the groups for which the exact value of their Noether number is given in [15]. For sake of completeness of the picture, we repeat that $\beta^{\mathbb{F}_{p}}(G)=\infty$ if $p$ divides $|G|$ by [43], whereas $\beta_{\mathrm{sep}}^{K}(G)\leq|G|$ for any $K$ (including the case when $\mathrm{char}(K)$ divides $|G|$).
@@ 3.3. Direct sum decompositions and multigradings
View the dual space $V_{i}^{*}$ as the subspace of $V^{*}$ consisting of the linear forms vanishing on $V_{j}$ for $j\neq i$. Then $V^{*}=\bigoplus_{i=1}^{k}V_{i}^{*}$. Choose a basis $x_{i}^{(j)}$ ($i=1,\dots,\dim(V_{j})$) in $V_{i}^{*}$, then our set of variables in $K[V]$ is $\{x_{i}^{(j)}\mid\ j=1,\dots,k;\ i=1,\dots,\dim(V_{j})\}$. Associated to the decomposition (8) there is an $\mathbb{N}_{0}^{k}$-grading on $K[V]$: the multihomogeneous component of multidegree $\alpha=(\alpha_{1},\dots,\alpha_{k})$ in $K[V]$ is spanned by the monomials having degree $\alpha_{j}$ in the variables belonging to $V_{j}^{*}$ for each $j=1,\dots,k$. Clearly, the $G$-action on $K[V]$ preserves this multigrading, and thus $K[V]^{G}$ is spanned by multihomogeneous elements. In the sequel whenever we mention a ‘multigrading’ not explicitly defined, this should refer to the multigrading that comes from a direct sum decomposition as in (8), which should be clear from the context. A typical way by which we shall exploit the multigrading is captured by the following obvious statement:
@@ 3.7. Relative invariants and the Hilbert ideal
For example, if $V$ is as in (11), then the variable $t_{\chi}$ is a relative invariant of weight $\chi$. Moreover, for a sequence $\chi^{(1)},\dots,\chi^{(k)}$ of elements of $\widehat{G}$, $t_{\chi^{(1)}}\cdots t_{\chi^{(k)}}$ is an indecomposable $G$-invariant if and only if $\chi^{(1)}\cdots\chi^{(k)}=1\in\widehat{G}$ and there is no $\ell<k$, $1\leq i_{1}<\cdots<i_{\ell}\leq k$ with $\chi^{(i_{1})}\cdots\chi^{(i_{\ell})}=1\in\widehat{G}$ (we say in this case that $\chi^{(1)},\dots,\chi^{(k)}$ is an irreducible product-one sequence over $\widehat{G}$, and we call $k$ its length). We say that $\chi^{(1)},\dots,\chi^{(k)}$ is a product-one free sequence over $\widehat{G}$ if there is no $1\leq i_{1}<\dots<i_{\ell}\leq k$ with $\chi^{(i_{1})}\cdots\chi^{(i_{\ell})}=1\in\widehat{G}$. The Davenport constant $\mathsf{D}(\widehat{G})$ of the finite abelian group $G$ is defined as the maximal length of an irreducible product-one sequence over $G$. Clearly, the maximal possible length of a product-one free sequence is $\mathsf{D}(\widehat{G})-1$. For more information about the Davenport constant and its relevance for the invariant theory of abelian groups see for example [14].
@@ 4. Groups with a cyclic subgroup of index two
Note that our representation is the direct sum of a $2$-dimensional and a $1$-dimensional representation. So by Lemma 3.5 it is sufficient to show that $v$ and $v^{\prime}$ can not be separated by a multihomogeneous invariant (cf. Section 3.3) of degree at most $m$. Suppose for contradiction that $f(v)\neq f(v^{\prime})$, where $f$ is a multihomogeneous invariant of degree at most $m$. So $f=t^{d}f_{1}(x_{1},x_{2})$. Since the projections of these two vectors in the $x_{1},x_{2}$ coordinate plane are the same, no invariant depending only on $x_{1}$ and $x_{2}$ can separate them. Thus we have $d>0$. The invariants depending only on $t$ are the polynomials in $t^{2}$, which take the same value (namely $1$) on these vectors, hence $\deg(f_{1})>0$. Since $t$ is $A$-invariant, $f_{1}(x_{1},x_{2})\in K[x_{1},x_{2}]^{A}$, so $f_{1}$ is a $K$-linear combination of $A$-invariant monomials in $x_{1}$ and $x_{2}$. The monomials having positive degree in $x_{2}$ vanish on both $v$ and $v^{\prime}$. It follows that $f_{1}$ must have an $A$-invariant monomial depending only on $x_{1}$. The only such monomials are the powers of $x_{1}^{m}$. Thus $\deg(f)\geq m+d>m$, a contradiction.
@@ 4.1. The groups $\mathrm{D}_{2n}\times\mathrm{C}_{2}$
implying in turn that $\beta_{\mathrm{sep}}^{K}(\mathrm{D}_{2n}\times\mathrm{C}_{2})\leq n+2$. The reverse inequality holds by Proposition 4.3. $\square$
@@ 5.2. The binary tetrahedral group
In this section $G:=\widetilde{\mathrm{A}}_{4}$ is the binary tetrahedral group: the special orthogonal group $\mathrm{SO}(3,\mathbb{R})$ has a subgroup isomorphic to $\mathrm{A}_{4}$ (the group of rotations mapping a regular tetrahedron into itself), and $G$ is isomorphic to the preimage of $\mathrm{A}_{4}$ under the classical two-fold covering $\mathrm{SU}(2,\mathbb{C})\to\mathrm{SO}(3,\mathbb{R})$. So $G$ can be embedded as a subgroup into the special unitary group $\mathrm{SU}(2,\mathbb{C})$, acting via its defining representation on $V:=\mathbb{C}^{2}$. Thus $V$ is a $G$-module, and it is well known that $\mathbb{C}[V]^{G}$ is generated by three homogeneous invariants $f_{6}$, $f_{8}$, $f_{12}$, having degree $6$, $8$, and $12$ (see [32, Lemma 4.1] or [42, Section 0.13]). Since $G$ (as a subgroup of $\mathrm{GL}(V)$) is not generated by pseudo-reflections, $\mathbb{C}[V]^{G}$ does not contain an algebraically independent separating set by [24], so the invariants $f_{6}$ and $f_{8}$ can not separate all the $G$-orbits in $V$. This implies that $\beta_{\mathrm{sep}}^{\mathbb{C}}(G,V)\geq 12$. A second proof valid for more general base fields is given below.
@@
having degrees $4,6,12,12$. In fact $h_{12}^{(1)}$ can be omitted, because we have $h_{12}^{(1)}=-4h_{6}^{2}$. Set $v:=[1,\xi^{2}]^{T}$ and $v^{\prime}:=\xi v$. Then $(x_{1}^{2}+x_{2}^{2})(v)=0=(x_{1}^{2}+x_{2}^{2})(v^{\prime})$, hence $f_{2}(v)=0=f_{2}(v^{\prime})$. It follows that $h_{6}(v)=0=h_{6}(v^{\prime})$. Moreover, $h_{8}(v)=16(\xi^{4})^{2}=16=16(\xi^{8})^{2}=h_{8}(v^{\prime})$. Thus no invariants of degree less than $12$ separate $v$ and $v^{\prime}$. However, they have different $G$-orbit, since $h_{12}^{(2)}(v)=4^{3}(\xi^{4})^{3}=-64$, whereas $h_{12}^{(2)}(v^{\prime})=4^{3}(\xi^{8})^{3}=64$. Thus we proved the inequality $\beta_{\mathrm{sep}}(G,V)\geq 12$, implying in turn $\beta_{\mathrm{sep}}^{K}(G)\geq 12$. On the other hand, by [11, Corollary 3.6] we know that $\beta^{K}(G)=12$, implying the reverse inequality $\beta_{\mathrm{sep}}^{K}(G)\leq 12$. ∎
@@ 5.6. The non-abelian group $(\mathrm{C}_{2}\times\mathrm{C}_{2})\rtimes\mathrm{C}_{4}$
has $c^{2}$ in its kernel and gives the defining representation of $\mathrm{D}_{8}$ as the group of symmetries of a square in the euclidean plane. The element $ab$ also belongs to the center of $G$, and the factor group $G/\langle ab\rangle\cong\mathrm{C}_{4}\times\mathrm{C}_{2}$ is generated by the coset of $c$ and the coset of $a$. Thus $G$ has the $1$-dimensional representations